OpenStudy (anonymous):

The perimeter of a rectangle is forty-eight feet. If the length measures sixteen feet, which is the width? A) 5 feet B) 8 feet C) ten feet D) sixteen feet

OpenStudy (anonymous):

i keep getting the wrong answer

OpenStudy (colouringme):

rectangle is consist of four side, 2 length and 2 width... it would only make sense to add each length of the side to find the perimeter. 2l + 2w = perimeter l x w = area The picture below shows you what you should measure...

OpenStudy (anonymous):

giving out medals to anyone who helps

OpenStudy (anonymous):

so 4

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

HELLO

OpenStudy (anonymous):

anyone

OpenStudy (colouringme):

I am so so sorry for this very late reply, OS was acting weird all day long, and now it's fine again. try to write down the equation, and watch out as you need to multiply the length and divide the width by two. Try to do it like this : 2l + 2w = Area 2x16 + 2w = 48 so the w is...